David and John's ages add up to twenty nine. Seven years ago David was twice the
age john. What are their present ages.

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

D+J=29
D-7=2(J-7)
so D=2J-14+7=2J-7
so 2J-7+J=29
3J=36
J=12
D=17
check 7 years ago they were 5 and 10.

Answer by greenestamps(13200)



You should know how to set up and solve the problem using formal algebra, as shown by the other tutor.

But solving a problem like this using logical reasoning and mental arithmetic is also good brain exercise.

If the sum of their ages is now 29, then 7 years ago the sum of their ages was 29-(2*7) = 29-14 = 15.

If 7 years ago David was twice the age of John, then with their two ages adding to 15 David was 10 and John was 5.

That was 7 years ago; so now their ages are 17 and 12.